OpenStudy (anonymous):

Solve (3a-2)=a+25/a+1

OpenStudy (bahrom7893):

multiply everything by a: a(3a-2) = a^2+25+a

OpenStudy (mertsj):

\[\frac{3a-2}{1}=\frac{a+25}{a+1}\]

OpenStudy (mertsj):

Cross multiply and solve.

OpenStudy (bahrom7893):

oh wait.. uhmm when are u people gonna learn how to use the parenthesis??

OpenStudy (mertsj):

Not going to happen bahrom

OpenStudy (bahrom7893):

come on merts... this is like the worst thing ever.. the other day someone asked a question, didn't post all the info given. Then after spending about 30 mins on the question, i decided to ask if there was any info given. She said yes and gave me two more pieces of info with which i could've solved the question in my head. I got annoyed and said.. couldn't you post everything instead of wasting both mine and your time.. and her response was... you don't have to answer this, it doesn't seem that u can do it anyway.. *smh*

OpenStudy (mertsj):

I know. It is very annoying but we cannot control the behavior of others.